OpenStudy (anonymous):

Choose the equation that represents the graph below: Graph of a line passing through points negative 2,0 and 0,-4 y = -2x - 4 y = 2x + 4 y = -2x + 4 y = 2x - 4

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

All of those options are in the form of \(y = mx + b\), where \(m\) is the slope, and \(b\) is the y-intercept. We have to find out which one passes through the points (-2, 0) and (0, -4). (0, -4) tells us that the y-intercept is -4. So that means that our equation has to have -4 at the end of it. \(y = mx - 4\) is what we have so far.

OpenStudy (igreen):

Okay, now lets find the slope between the two points by plugging them into the slope formula. \(m = \dfrac{y_2-y_1}{x_2-x_1}\) \(m = \dfrac{-4-0}{~~~0+2}\) Add & Subtract: \(m = \dfrac{-4}{~~~2}\) Can you divide -4 / 2?

OpenStudy (anonymous):

-2

OpenStudy (igreen):

Yep! So that gives us \(y = -2x - 4\).